OpenStudy (anonymous):

Please help! Algebra assigment, only need one question Satisfied with his notes, the guardian lets you pass into the next chamber. As you enter the dimly lit room, McMerlock points out that the floor tiles have numbers on them, and that some floor tiles are missing. On the ceiling is painted a cryptic message. 3. “Only the solutions of g(x) will lead you safely across.” Find the solutions of g(x). Show each step. my function is g(x) = x^2 =2

OpenStudy (anonymous):

I'm really bad with solving functions

OpenStudy (loser66):

g(x) = x^2+ 2, I assume, right/

OpenStudy (anonymous):

whoops my bad, I mean't a plus sign!

OpenStudy (loser66):

so, you will end up at complex roots?

OpenStudy (anonymous):

two complex solutions

OpenStudy (loser66):



OpenStudy (anonymous):

is that the solution

OpenStudy (anonymous):

@AllTehMaffs Can you help I don't exactly understand what they were saying

OpenStudy (anonymous):

So if you have your function \[g(x) = x^2+2\] then to find the solutions, you set the equation equal to zero \[g(x)=x^2+2=0\] Which is of the form \[f(x)=ax^2+bx+c=0\] where with g(x) \[a=1 \\ b=0 \\ c=2\] And then use the quadratic formula to find the solutions \[x=\frac{-b±\sqrt{b^2-4ac}}{2a} = \frac{±\sqrt{-4*2}}{2} = \frac{±2i\sqrt{2}}{2}=±i\sqrt{2}\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

^_^